OpenStudy (anonymous):

A rectangular painting has a diagonal measure of 13 inches and an area of 60 square inches. Use the formula for the area of a rectangle and the Pythagorean Theorem to find the length and width of the painting

OpenStudy (anonymous):

This question you can answer using "special triangles." The most common of these triangles is obviously the 3 4 5 triangle. But another of these is the 5,12,13 triangle. Since you know the hypotenuse is 13 you can try 5 and 12 as the length and width. Luckily, 5*12=60! Using the pythagorean theorem and the area of a rectangle is a little more difficult. you have to set up a system of equations. You know: x^2 + y^2 = (13)^2 and x*y=60. Using these two equations, you should be able to find the values of x and y

OpenStudy (anonymous):

thank you I realize my mistake i was using the perimeter 2l + 2w

OpenStudy (anonymous):

oops! my pleasure. good luck!